Find standard deviation for the following data:

X 6 9 12 15 18

f 7 12 13 10 8


1
Expert's answer
2021-01-15T15:06:36-0500

To find the standard deviation let us take "A"( assumed mean) "=12" . Then calculate the following table such as




Now the formula for standard deviation is "\sigma =\sqrt{\frac{\Sigma f.d^2}{N}-(\frac{\Sigma f.d}{N})^2}" where "N=\Sigma f"

From the above table we get , "\sigma =\sqrt{\frac{738}{50}-(\frac {0}{50})^2}" "=\sqrt{14.76}\approx 3.84" (correct upto two decimal places)

Therefore required standard deviation is "=3.84" (approximately)